Identify how each situation will affect consumption (C) and Saving (S) in an economy. The Federal...

  1. Identify how each situation will affect consumption (C) and Saving (S) in an economy.
    1. The Federal Reserve reduces key interest rates.
    2. The government enacts a cut to individual income taxes.
    3. You are expecting to lose your job next month.
    4. The value of household debts increases.
  1. Currently the value of Real GDP in an economy is $12 trillion and the level of Government spending (G) is $2.5 trillion. We observe that when Government Spending increases to $3 trillion, the value of Real GDP increases to $13.4 trillion. What is the value of the Multiplier for this economy?

  1. By how much will Real GDP change if firms increase their investments by $8 billion and the MPC = 0.7? If the MPC = 0.65?
Answer #1

a)

When Fed reduces key interest rate. it will increase money supply and interest rate will decline in market. So , Cost of holding money declines, so Consumption will rise and saving will decline.

b)

Fall in income tax would increase consumer income, so both consumption and savings are likely to rise.

c)

Lose of job means no income next month onwards, so consumption will fall and consumer will seek to increase saving to compensate for jobless period.

d)

When value of household debt rises, consumer will reduce the consumption and would raise saving to payoff debt.

Multiplier = Change in GDP / Change in Expenditure

= 1.4 /0.5

= 2.8

GDP increase when MPC = 0.7

Multiplier = 1/ - 0.7

= 1/0.3

= 3.33

Rise in GDP = 8*3.33

= 26.64

GDP increase in when MPC = 0.65

Multiplier = 1/1-0.65

= 2.85

Rise in GDP = 2.85*8

= 22.8